40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1664     Entry time: Wed Jul 13 00:06:17 2016
Author: Antonio, Andrew 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: North path construction progress 3 

Summary
Following ID 1649 & 1663 electronics for the FSS of the North path has been hooked up.
The North cavity was ready for locking today.

Conclusion
The error-signal of the North PDH does not look right. The cavity can be locked but it is
not an easy operation for now as it was before. We need to investigate on the error-signal
and nail down the reason.



Notes:

  • Initially the err-sig had a weird shape and a low SNR (~2), later it became for unknown reasons a proper
    PDH with still the same SNR.
  • Attempts to lock the North cavity have been done anyway. Without driving the BBEOM 
    cavity stayied locked for half an hour. 
  • The error signal was taken at the monitor of the TTFSS and at the FSS box on the table (out1 and out 2); they were similar

 

 

ELOG V3.1.3-